# Data-driven algorithm design using neural networks with applications to branch-and-cut

### v1

@@ Abstract
Data-driven algorithm design is a paradigm that uses statistical and machine learning techniques to select from a class of algorithms for a computational problem an algorithm that has the best expected performance with respect to some (unknown) distribution on the instances of the problem. We build upon recent work in this line of research by introducing the idea where, instead of selecting a single algorithm that has the best performance, we allow the possibility of selecting an algorithm based on the instance to be solved. In particular, given a representative sample of instances, we learn a neural network that maps an instance of the problem to the most appropriate algorithm for that instance. We formalize this idea and derive rigorous sample complexity bounds for this learning problem, in the spirit of recent work in data-driven algorithm design. We then apply this approach to the problem of making good decisions in the branch-and-cut framework for mixed-integer optimization (e.g., which cut to add?). In other words, the neural network will take as input a mixed-integer optimization instance and output a decision that will result in a small branch-and-cut tree for that instance. Our computational results provide evidence that our particular way of using neural networks for cut selection can make a significant impact in reducing branch-and-cut tree sizes, compared to previous data-driven approaches.
Keywords Integer programming, data-driven algorithm design, branch-and-cut, branch-and-bound, sample complexity, neural networks, learning theory
@@ 1 Background and motivation
Often there are several competing algorithms for a computational problem and no single algorithm dominates all the others. The choice of an algorithm in such cases is often dictated by the “typical" instance one expects to see, which may differ from one application context to another. Data-driven algorithm design has emerged in recent years as a way to formalize this question of algorithm selection and draws upon statistical and machine learning techniques; see (Balcan, 2020) for a survey and references therein. More formally, suppose one has a class $\mathcal{I}$ of instances of some computational problem with some unknown distribution, and class of algorithms that can solve this problem parameterized by some “tunable parameters". Suppose that for each setting of the parameters, the corresponding algorithm can be evaluated by a score function that tells us how well the algorithm does on different instances (e.g., running time, memory usage etc.). We wish to find the set of parameters that minimizes (or maximizes, depending on the nature of the score) the expected score on the instances (with respect to the unknown distribution), after getting access to an i.i.d. sample of instances from the distribution. For instance, $\mathcal{I}$ could be a family of mixed-integer optimization problems and the class of algorithms are branch-and-cut methods with their different possible parameter settings, and the score function could be the size of the branch-and-cut tree.
@@
In (Balcan et al., 2021a), the authors prove a central result on the sample complexity of this problem: if for any fixed instance in $\mathcal{I}$, the score function has a piecewise structure as a function of the parameters, where the number of pieces is upper bounded by a constant $R$ independent of the instance, the pieces come from a family of functions of “low complexity", and the regions partitioning the parameter space into pieces have “low complexity" in terms of their shape, then the sample complexity is a precise function of $R$ and these two complexity numbers (formalized by the notion of pseudo-dimension – see below). See Theorem 3.3 in (Balcan et al., 2021a) for details. The authors then proceed to apply this key result to a diverse variety of algorithmic problems ranging from computational biology, computational economics to integer optimization.
@@
Our work in this paper is motivated by the following consideration. In many of these applications, one would like to select not a single algorithm, i.e., a single setting of the parameters, but would like to decide on the parameters after receiving a new instance of the problem. In other words, we would like to learn not the best parameter, but the best mapping from instances to parameters. We consider this version of the problem, where the mapping is taken from a class of neural network functions. At a high level, switching the problem from finding the “best" parameter to finding the “best" mapping in a parameterized family of mappings gives the same problem: we now have a new set of parameters – the ones parameterizing the mappings (neural networks) – and we have to select the best parameter. And indeed, the result from (Balcan et al., 2021a) can be applied to this new setting, as long as one can prove good bounds on pseudo-dimension in the space of these new parameters. The key point of our result is two fold:
@@
    Even if original space of parameters for the algorithms is amenable to the analysis of pseudo-dimension as done in (Balcan et al., 2021a), it is not clear that this immediately translates to a similar analysis in the parameter space of the neural networks, because the “low complexity piecewise structure" in the original parameter space may not be preserved in the neural parameter space. Our proof technique, while very much inspired by the conceptual insights in (Balcan et al., 2021a), has to solve the problem ab initio.
@@
    We suspect that our proof technique results in smaller sample complexity bounds, compared to what one would obtain if one could do the analysis from (Balcan et al., 2021a) in the parameter space of the neural network directly. However, we do not have quantitative evidence of this yet because this analysis seems difficult to carry out in the neural parameter space.
Our approach is also related to recent work on algorithm design with predictions; see (Mitzenmacher and Vassilvitskii, 2022) for a short introduction and the references therein for more thorough surveys and recent work. While one can interpret that line of work as also mapping instances to the best parameters in a specific algorithmic framework, our emphasis and the nature of our results is quite different from the focus in that area.
Our main sample complexity results for this neural approach to data-driven algorithm design is presented in Section 2. We use these results as a theoretical basis for making progress on the fundamental problem of cut selection in mixed-integer linear optimization, using neural networks. We turn to this discussion next.
@@ 1.1 Applications in branch-and-cut methods for mixed-integer linear optimization
Mixed-integer linear optimization is a powerful tool that is used in a diverse number of application domains. Branch-and-cut is the solution methodology of choice for all state-of-the-art solvers for mixed-integer optimization that is founded upon a well-developed theory of convexification and systematic enumeration (Schrijver, 1986; Nemhauser and Wolsey, 1988; Conforti et al., 2014). However, even after decades of theoretical and computational advances, several key aspects of branch-and-cut are not well understood. During the execution of the branch-and-cut algorithm on an instance, the algorithm has to repeatedly make certain decisions such as which node of the search tree to process next, whether one should branch or add cutting planes, which cutting planes to add, or which branching strategy to use. The choices that give a small tree size for a particular instance may not be good choices for a different instance and result in a much larger tree. Thus, adapting these choices to the particular instance can be beneficial for overall efficiency. Of course, the strategies already in place in the best implementations of branch-and-cut have to adapt to the instances. For example, certain cutting plane choices may not be possible for certain instances. But even beyond that, there are certain heuristics in place that adapt these choices to the instance. These heuristics have been arrived at by decades of computational experience from practitioners. The goal in recent years is to provide a data driven approach to making these choices. In a recent series of papers  (Balcan et al., 2021a, b, c, 2022, 2018), the authors apply the general sample complexity result from (Balcan et al., 2021a) in the specific context of branch-and-cut methods for mixed-integer linear optimization to obtain several remarkable and first-of-their-kind results. We summarize here those results that are most relevant to the cut selection problem since this is the focus of our work.
@@
    In (Balcan et al., 2021b), the authors consider the problem of selecting the best Chvátal-Gomory (CG) cutting plane (or collection of cutting planes) to be added at the root node of the branch-and-cut tree. Thus, the “tunable parameters" are the possible multipliers to be used to generate the CG cut at the root node. The score function is the size of the resulting branch-and-cut tree. The authors observe that for a fixed instance of a mixed-integer linear optimization problem, there are only finitely many CG cuts possible and the number can be bounded explicitly in terms of entries of the linear constraints of the problem. Via a sophisticated piece of analysis, this gives the required piecewise structure on the space of multipliers to apply the general result explained above. This gives concrete sample complexity bounds for choosing the multipliers with the best expected performance. See Theorem 3.3, 3.5 and 3.6 in (Balcan et al., 2021b) for details. Note that this result is about selecting a single set of multipliers/cutting planes that has the best expected performance across all instances. This contrasts with selecting a good strategy to select multipliers depending on the instance, that has good expected performance (see point 2. below).
@@
    The authors in (Balcan et al., 2021b) also consider the problem of learning a good strategy that selects multipliers based on the instance. In particular, they consider various auxiliary score functions used in integer optimization practice, that map a pair of instance $I$ and a cutting plane $c$ for $I$ that measures how well $c$ will perform for processing $I$. The strategy will be a linear function of these auxiliary scores, i.e., a weighted average of these auxiliary scores, and the learning problem becomes the problem of finding the best linear coefficients for the different auxiliary scores. So now these linear coefficients become the “tunable parameters" for the general result from (Balcan et al., 2021a). It is not hard to find the piecewise structure in the space of these new parameters, given the analysis in the space of CG cut multipliers from point 1. above. This then gives concrete sample complexity bounds for learning the best set of weights for these auxiliary score functions for cut selection. See Theorem 4.1 and Corollary 4.2 in (Balcan et al., 2021b) for details.
@@
    In all the previous results discussed above, the cutting planes considered were CG cuts, or it was essentially assumed that there are only a finite number of cutting planes available at any stage of the branch-and-cut algorithm. In the most recent paper (Balcan et al., 2022), the authors consider general cutting plane paradigms, and also consider the possibility of allowing more general strategies to select cutting planes beyond using weighted combinations of auxiliary score functions. They uncover the subtlety that allowing general mappings from instances to cutting planes can lead to infinite sample complexity and learning such mappings could be impossible, if the class of mappings is allowed to be too rich. See Theorem 5.1 in (Balcan et al., 2022). This point will be important when we discuss our approach below.
    On the positive side, they show that the well-known class of Gomory-Mixed-Integer (GMI) cuts has a similar structure to CG cuts, and therefore, using similar techniques as discussed above, they derive sample complexity bounds for selecting GMI cuts at the root node. See Theorem 5.5 in  (Balcan et al., 2022). As far as we understand, the analysis should extend to the problem of learning weighted combinations of auxiliary score functions to select the GMI cuts as well using the same techniques as (Balcan et al., 2021b), although the authors do not explicitly do this in (Balcan et al., 2022).
@@
As already discussed, our theoretical sample complexity work in cut selection is closest in spirit to the work in (Balcan et al., 2021a, b, c, 2022, 2018). However, there are several other works in the literature that use machine learning ideas to approach the problem of cut selection (Deza and Khalil, 2023). Tang et al. (2020) initiated the exploration of applying Reinforcement Learning (RL) to select CG cuts derived from the simplex tableau. Huang et al. (2022) apply supervised learning to rank a “bag of cuts” from a set of cuts to reduce the total runtime. More recently, the authors in (Turner et al., 2023) propose a novel strategy that use RL and Graph Neural Networks (GNN) to select instance-dependent weights for the combination of auxiliary score functions.
Our investigation, as detailed in Section 4, distinguishes itself from these prior explorations in several key aspects:
@@
    Many prior studies aimed at improving the objective value rather than directly reducing the branch-and-cut runtime—with the exception of Huang et al. (2022), who explored this through supervised learning. To the best of our knowledge, our RL-based model is the first to directly target the reduction of the branch-and-cut tree size as its reward metric, which is strongly correlated with the overall running time.
@@
    Prior deep learning approaches are not underpinned by theoretical guarantees, such as sample complexity bounds. Our empirical work takes the theoretical insights for the branch-and-cut problem presented in Theorem 2.3 and Theorem 3.3 as its basis.
@@ 2.1.1 Background from learning theory
$2^{t}=\left|\left\{\left(\operatorname{sgn}(f(I_{1})-s^{1}),\dots,\operatorname{sgn}(f(I_{t})-s^{t})\right):f\in\mathcal{F}\right\}\right|.$
@@
The following is a fundamental result in empirical processes theory and is foundational for the above learning problem; see, e.g., Chapters 17, 18 and 19 in (Anthony et al., 1999), especially Theorem 19.2.
@@ 2.2 Our results
In this study, we extend the framework introduced by Balcan et al. (2021a) to explore the learnability of tunable algorithmic parameters through neural networks. Consider a computational problem given by a family of instances $\mathcal{I}$. Let us say we have a suite of algorithms for this problem, parameterized by parameters in $\mathcal{P}$. We also have a score function that evaluates how well a particular algorithm, given by specific settings of the parameters, performs on a particular instance. In other words, the score function is given by $S:\mathcal{I}\times\mathcal{P}\rightarrow[0,B]$, where $B\in\mathbb{R}_{+}$ determines a priori upper bound on the score. The main goal of data-driven algorithm design is to find a particular algorithm in our parameterized family of algorithms – equivalently, find a parameter setting $\mathbf{p}\in\mathcal{P}$ – that minimizes the expected score on the family of instances with respect to an unknown distribution on $\mathcal{I}$, given access to a sample of i.i.d instances from the distribution. This then becomes a special case of the general learning problem (1), where $h=S$ and one can provide precise sample complexity bounds via Theorem 2.3, if one can bound the pseudo-dimension of the corresponding hypothesis class. A bound on this pseudo-dimension is precisely the central result in (Balcan et al., 2021a); see the discussion in Section 1.
@@
1. An encoder function $\operatorname{Enc}:\mathcal{I}\rightarrow\mathbb{R}^{d}$ is defined to convert an instance $I\in\mathcal{I}$ into a vector $\mathbf{x}=\operatorname{Enc}(I)\in\mathbb{R}^{d}$, facilitating the instances to be suitably processed by a neural network. A simple example of such an encoder could be a compilation of all the instance’s numerical data into a single vector; but one can allow encodings that use some predetermined features of the instances.
2. A family of neural network mappings, denoted as $N^{\sigma}:\mathbb{R}^{d}\times\mathbb{R}^{W}\rightarrow\mathbb{R}^{\ell}$, is utilized. These mappings are characterized by an activation function $\sigma$, and a fixed architecture represented by $\bm{w}=[d,w_{1},\ldots,w_{L},\ell]\in\mathbb{Z}_{+}^{L+2}$. For any given neural network parameters $\mathbf{w}\in\mathbb{R}^{W}$, this network maps an encoded instance $\mathbf{x}\in\mathbb{R}^{d}$ into $\mathbf{y}:=N^{\sigma}(\mathbf{x},\mathbf{w})\in\mathbb{R}^{\ell}$.
3. A squeezing activation function, $\sigma^{\prime}:\mathbb{R}\rightarrow[0,1]$, is introduced to adjust the neural network’s output to the parameter space $\mathcal{P}$. The parameter $\mathbf{p}\in\mathcal{P}$ is computed by $\mathbf{p}_{i}=\eta_{i}+(\tau_{i}-\eta_{i})\sigma^{\prime}(\mathbf{y}_{i})$ for $i=1,\ldots,\ell$.
@@
Our first result employs linear threshold neural networks for generating algorithm parameters, inspired by their analytically tractable structure and rich expressive capabilities, as supported by findings in (Khalife et al., 2023).
@@
Under the same conditions as Theorem 2.5 and 2.6, but with $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{r}\}$, we have
$\operatorname{Pdim}(\mathcal{F}^{S}_{N^{\operatorname{sgn}},\operatorname{Sigmoid}})$ $=\mathcal{O}\left(W\log(Ur)\right),$
$\operatorname{Pdim}(\mathcal{F}^{S}_{N^{\operatorname{ReLU}},\operatorname{CReLU}})$ $=\mathcal{O}\left(LW\log(U+r)\right).$

### v2

# Sample Complexity of Algorithm Selection Using Neural Networks and Its Applications to Branch-and-Cut

@@ Abstract
Data-driven algorithm design is a paradigm that uses statistical and machine learning techniques to select from a class of algorithms for a computational problem an algorithm that has the best expected performance with respect to some (unknown) distribution on the instances of the problem. We build upon recent work in this line of research by considering the setup where, instead of selecting a single algorithm that has the best performance, we allow the possibility of selecting an algorithm based on the instance to be solved, using neural networks. In particular, given a representative sample of instances, we learn a neural network that maps an instance of the problem to the most appropriate algorithm for that instance. We formalize this idea and derive rigorous sample complexity bounds for this learning problem, in the spirit of recent work in data-driven algorithm design. We then apply this approach to the problem of making good decisions in the branch-and-cut framework for mixed-integer optimization (e.g., which cut to add?). In other words, the neural network will take as input a mixed-integer optimization instance and output a decision that will result in a small branch-and-cut tree for that instance. Our computational results provide evidence that our particular way of using neural networks for cut selection can make a significant impact in reducing branch-and-cut tree sizes, compared to previous data-driven approaches.
@@ 1 Background and motivation
Often there are several competing algorithms for a computational problem and no single algorithm dominates all the others. The choice of an algorithm in such cases is often dictated by the “typical" instance one expects to see, which may differ from one application context to another. Data-driven algorithm design has emerged in recent years as a way to formalize this question of algorithm selection and draws upon statistical and machine learning techniques; see Balcan (2020) for a survey and references therein. More formally, suppose one has a class $\mathcal{I}$ of instances of some computational problem with some unknown distribution, and class of algorithms that can solve this problem parameterized by some “tunable parameters". Suppose that for each setting of the parameters, the corresponding algorithm can be evaluated by a score function that tells us how well the algorithm does on different instances (e.g., running time, memory usage etc.). We wish to find the set of parameters that minimizes (or maximizes, depending on the nature of the score) the expected score on the instances (with respect to the unknown distribution), after getting access to an i.i.d. sample of instances from the distribution. For example, $\mathcal{I}$ could be a family of mixed-integer optimization problems and the class of algorithms are branch-and-cut methods with their different possible parameter settings, and the score function could be the size of the branch-and-cut tree.
@@
In Balcan et al. (2021a), the authors prove a central result on the sample complexity of this problem: if for any fixed instance in $\mathcal{I}$, the score function has a piecewise structure as a function of the parameters, where the number of pieces is upper bounded by a constant $R$ independent of the instance, the pieces come from a family of functions of “low complexity", and the regions partitioning the parameter space into pieces have “low complexity" in terms of their shape, then the sample complexity is a precise function of $R$ and these two complexity numbers (formalized by the notion of pseudo-dimension – see below). See Theorem 3.3 in Balcan et al. (2021a) for details. The authors then proceed to apply this key result to a diverse variety of algorithmic problems ranging from computational biology, computational economics to integer optimization.
@@
Our work in this paper is motivated by the following consideration. In many of these applications, one would like to select not a single algorithm, i.e., a single setting of the parameters, but would like to decide on the parameters after receiving a new instance of the problem. In other words, we would like to learn not the best parameter, but the best mapping from instances to parameters. We consider this version of the problem, where the mapping is taken from a class of neural network functions. At a high level, switching the problem from finding the “best” parameter to finding the “best” mapping in a parameterized family of mappings gives the same problem: we now have a new set of parameters – the ones parameterizing the mappings (neural networks) – and we have to select the best parameter. And indeed, the result from Balcan et al. (2021a) can be applied to this new setting, as long as one can prove good bounds on pseudo-dimension in the space of these new parameters. The key point of our result is two fold:
@@
    Even if original space of parameters for the algorithms is amenable to the analysis of pseudo-dimension as done in Balcan et al. (2021a), it is not clear that this immediately translates to a similar analysis in the parameter space of the neural networks, because the “low complexity piecewise structure" from the original parameter space may not be preserved.
@@
    We suspect that our proof technique results in tighter sample complexity bounds, compared to what one would obtain if one could do the analysis from Balcan et al. (2021a) in the parameter space of the neural network directly. However, we do not have quantitative evidence of this yet because their analysis seems difficult to carry out directly in the neural parameter space.
One of the foundational works in the field of selecting algorithms based on specific instances is Rice (1976), and recently, Gupta and Roughgarden (2016); Balcan et al. (2021c) have explored the sample complexity of learning mappings from instances to algorithms for particular problems. Our approach is also related to recent work on algorithm design with predictions; see Mitzenmacher and Vassilvitskii (2022) for a short introduction and the references therein for more thorough surveys and recent work. However, our emphasis and the nature of our results are quite different from the focus of previous research. We establish sample complexity bounds for a general learning framework that employs neural networks to map instances to algorithms, which is suitable for handling some highly complex score functions, such as the size of the branch-and-cut tree. The introduction of neural networks in our approach bring more intricate technical challenges than traditional settings like linear predictors or regression trees Gupta and Roughgarden (2016). On the application side, as long as we know some information about the algorithms (see Theorems 2.5 and 2.6), our results can be directly applied. This is demonstrated in this paper through applications to various cutting plane selection problems in branch-and-cut in Propositions 3.33.53.6 and 3.7. Several references to fundamental applied work in data-driven algorithm selection can be found in the Balcan (2020); Gupta and Roughgarden (2016); Balcan et al. (2021a, c).
@@ 1.1 Applications in branch-and-cut methods for mixed-integer linear optimization
Mixed-integer linear optimization is a powerful tool that is used in a diverse number of application domains. Branch-and-cut is the solution methodology of choice for all state-of-the-art solvers for mixed-integer optimization that is founded upon a well-developed theory of convexification and systematic enumeration Schrijver (1986); Nemhauser and Wolsey (1988); Conforti et al. (2014). However, even after decades of theoretical and computational advances, several key aspects of branch-and-cut are not well understood. During the execution of the branch-and-cut algorithm on an instance, the algorithm has to repeatedly make certain decisions such as which node of the search tree to process next, whether one should branch or add cutting planes, which cutting planes to add, or which branching strategy to use. The choices that give a small tree size for a particular instance may not be good choices for a different instance and result in a much larger tree. Thus, adapting these choices to the particular instance can be beneficial for overall efficiency. Of course, the strategies already in place in the best implementations of branch-and-cut have to adapt to the instances. For example, certain cutting plane choices may not be possible for certain instances. But even beyond that, there are certain heuristics in place that adapt these choices to the instance. These heuristics have been arrived at by decades of computational experience from practitioners. The goal in recent years is to provide a data driven approach to making these choices. In a recent series of papers  Balcan et al. (2021a, d, b, 2022, 2018), the authors apply the general sample complexity result from Balcan et al. (2021a) in the specific context of branch-and-cut methods for mixed-integer linear optimization to obtain several remarkable and first-of-their-kind results. We summarize here those results that are most relevant to the cut selection problem since this is the focus of our work.
@@
    In Balcan et al. (2021d), the authors consider the problem of selecting the best Chvátal-Gomory (CG) cutting plane (or collection of cutting planes) to be added at the root node of the branch-and-cut tree. Thus, the “tunable parameters" are the possible multipliers to be used to generate the CG cut at the root node. The score function is the size of the resulting branch-and-cut tree. The authors observe that for a fixed instance of a mixed-integer linear optimization problem, there are only finitely many CG cuts possible and the number can be bounded explicitly in terms of entries of the linear constraints of the problem. Via a sophisticated piece of analysis, this gives the required piecewise structure on the space of multipliers to apply the general result explained above. This gives concrete sample complexity bounds for choosing the multipliers with the best expected performance. See Theorem 3.3, 3.5 and 3.6 in Balcan et al. (2021d) for details. Note that this result is about selecting a single set of multipliers/cutting planes that has the best expected performance across all instances. This contrasts with selecting a good strategy to select multipliers depending on the instance, that has good expected performance (see point 2. below).
@@
    The authors in Balcan et al. (2021d) also consider the problem of learning a good strategy that selects multipliers based on the instance. In particular, they consider various auxiliary score functions used in integer optimization practice, that map a pair of instance $I$ and a cutting plane $c$ for $I$ that measures how well $c$ will perform for processing $I$. The strategy will be a linear function of these auxiliary scores, i.e., a weighted average of these auxiliary scores, and the learning problem becomes the problem of finding the best linear coefficients for the different auxiliary scores. So now these linear coefficients become the “tunable parameters" for the general result from Balcan et al. (2021a). It is not hard to find the piecewise structure in the space of these new parameters, given the analysis in the space of CG cut multipliers from point 1. above. This then gives concrete sample complexity bounds for learning the best set of weights for these auxiliary score functions for cut selection. See Theorem 4.1 and Corollary 4.2 in Balcan et al. (2021d) for details.
@@
    In all the previous results discussed above, the cutting planes considered were CG cuts, or it was essentially assumed that there are only a finite number of cutting planes available at any stage of the branch-and-cut algorithm. In the most recent paper Balcan et al. (2022), the authors consider general cutting plane paradigms, and also consider the possibility of allowing more general strategies to select cutting planes beyond using weighted combinations of auxiliary score functions. They uncover the subtlety that allowing general mappings from instances to cutting planes can lead to infinite sample complexity and learning such mappings could be impossible, if the class of mappings is allowed to be too rich. See Theorem 5.1 in Balcan et al. (2022). This point will be important when we discuss our approach below.
    On the positive side, they show that the well-known class of Gomory-Mixed-Integer (GMI) cuts has a similar structure to CG cuts, and therefore, using similar techniques as discussed above, they derive sample complexity bounds for selecting GMI cuts at the root node. See Theorem 5.5 in  Balcan et al. (2022). As far as we understand, the analysis should extend to the problem of learning weighted combinations of auxiliary score functions to select the GMI cuts as well using the same techniques as Balcan et al. (2021d), although the authors do not explicitly do this in Balcan et al. (2022).
@@
As already discussed above, our theoretical sample complexity work in cut selection is closest in spirit to the work in Balcan et al. (2021a, d, b, 2022, 2018). However, there are several other works in the literature that use machine learning ideas to approach the problem of cut selection; see Deza and Khalil (2023) for a survey. Tang et al. Tang et al. (2020) initiated the exploration of applying Reinforcement Learning (RL) to select CG cuts derived from the simplex tableau. Huang et al. Huang et al. (2022) apply supervised learning to rank a “bag of cuts” from a set of cuts to reduce the total runtime. More recently, the authors in Turner et al. (2023) propose a novel strategy that use RL and Graph Neural Networks (GNN) to select instance-dependent weights for the combination of auxiliary score functions.
Our computational investigations, as detailed in Section 4, distinguishes itself from these prior computational explorations in several key aspects:
@@
    Many prior studies aimed at improving the objective value rather than directly reducing the branch-and-cut runtime—with the exception of Huang et al. Huang et al. (2022), who explored this through supervised learning. To the best of our knowledge, our RL-based model is the first to directly target the reduction of the branch-and-cut tree size as its reward metric, which is strongly correlated with the overall running time.
@@
    Prior deep learning approaches are not underpinned by theoretical guarantees, such as sample complexity bounds. Our empirical work takes the theoretical insights for the branch-and-cut problem presented in Theorem 2.3 and Proposition 3.3 as its basis.
The limitations of our approach are discussed in Section 5.
@@ 2.1.1 Background from learning theory
$2^{t}=\left|\left\{\left(\operatorname{sgn}(f(I_{1})-s_{1}),\dots,\operatorname{sgn}(f(I_{t})-s_{t})\right):f\in\mathcal{F}\right\}\right|.$
@@
The following is a fundamental result in empirical processes theory and is foundational for the above learning problem; see, e.g., Chapters 17, 18 and 19 in Anthony et al. (1999), especially Theorem 19.2.
@@ 2.2 Our results
In this study, we extend the framework introduced by Balcan et al. Balcan et al. (2021a) to explore the learnability of tunable algorithmic parameters through neural networks. Consider a computational problem given by a family of instances $\mathcal{I}$. Let us say we have a suite of algorithms for this problem, parameterized by parameters in $\mathcal{P}$. We also have a score function that evaluates how well a particular algorithm, given by specific settings of the parameters, performs on a particular instance. In other words, the score function is given by $S:\mathcal{I}\times\mathcal{P}\rightarrow[0,B]$, where $B\in\mathbb{R}_{+}$ determines a priori upper bound on the score. The main goal of data-driven algorithm design is to find a particular algorithm in our parameterized family of algorithms – equivalently, find a parameter setting $\mathbf{p}\in\mathcal{P}$ – that minimizes the expected score on the family of instances with respect to an unknown distribution on $\mathcal{I}$, given access to a sample of i.i.d instances from the distribution. This then becomes a special case of the general learning problem (1), where $h=S$ and one can provide precise sample complexity bounds via Theorem 2.3, if one can bound the pseudo-dimension of the corresponding hypothesis class. A bound on this pseudo-dimension is precisely the central result in Balcan et al. (2021a); see the discussion in Section 1.
@@
  1. 1.
    An encoder function $\operatorname{Enc}:\mathcal{I}\rightarrow\mathbb{R}^{d}$ is defined to convert an instance $I\in\mathcal{I}$ into a vector $\mathbf{x}=\operatorname{Enc}(I)\in\mathbb{R}^{d}$, facilitating the instances to be suitably processed by a neural network. A simple example of such an encoder could be a compilation of all the instance’s numerical data into a single vector; but one can allow encodings that use some predetermined features of the instances.
  2. 2.
    A family of neural network mappings, denoted as $N^{\sigma}:\mathbb{R}^{d}\times\mathbb{R}^{W}\rightarrow\mathbb{R}^{\ell}$, is utilized. These mappings are characterized by an activation function $\sigma$, and a fixed architecture represented by $\bm{w}=[d,w_{1},\ldots,w_{L},\ell]\in\mathbb{Z}_{+}^{L+2}$. For any given neural network parameters $\mathbf{w}\in\mathbb{R}^{W}$, this network maps an encoded instance $\mathbf{x}\in\mathbb{R}^{d}$ into $\mathbf{y}:=N^{\sigma}(\mathbf{x},\mathbf{w})\in\mathbb{R}^{\ell}$.
  3. 3.
    A squeezing activation function, $\sigma^{\prime}:\mathbb{R}\rightarrow[0,1]$, is introduced to adjust the neural network’s output to the parameter space $\mathcal{P}$. The parameter $\mathbf{p}\in\mathcal{P}$ is computed by $\mathbf{p}_{i}=\eta_{i}+(\tau_{i}-\eta_{i})\sigma^{\prime}(\mathbf{y}_{i})$ for $i=1,\ldots,\ell$.
@@
Our first result employs linear threshold neural networks for generating algorithm parameters, inspired by their analytically tractable structure and rich expressive capabilities, as supported by findings in Khalife et al. (2023).
@@
Under the same conditions as Theorem 2.5 and 2.6, but with $\mathcal{P}=\{\mathbf{p}_{1},\dots,\mathbf{p}_{r}\}$,
$\operatorname{Pdim}(\mathcal{F}^{S}_{N^{\operatorname{sgn}},\operatorname{Sigmoid}})=\mathcal{O}\left(W\log(Ur)\right)\text{ and }\operatorname{Pdim}(\mathcal{F}^{S}_{N^{\operatorname{ReLU}},\operatorname{CReLU}})=\mathcal{O}\left(LW\log(U+r)\right).$